OpenStudy (anonymous):

Clive loaned his brother some money to buy a new truck, his brother agreed to repay the loan in 1 year with 3% interest. Clive earned $75 interest on the loan. How much did clive loan his brother?

OpenStudy (anonymous):

If $75=3%, what is 100%? Using cross-multiplying, 75/3=x/100 3x=7500 x=2500 So, Clive loaned his brother $2,500.